1. Jerry is 13 In five years his age will be 3/2 the age of his sister Maria How old will Maria be in 3 years

Answer by ewatrrr(24785)


Let x be Maria's age now
In 5 years...
13+ 5 = (3/2)(x + 5)
18 = (3/2)x + 15/2
Multiply thru by 2
36 = 3x + 15
21 = 3x
x = 7, Maria's age NOW. Maria will be 10 in 3 years